40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 21 19:28:03 2021, Anchal, Update, PSL, PSL/IFO recovery 
    Reply  Wed Apr 21 19:43:20 2021, Koji, Update, PSL, New HEPA speed control P_20210421_193637.jpgP_20210421_193627.jpg
       Reply  Thu Apr 22 14:41:55 2021, Chub, Update, PSL, New HEPA speed control 40M_PSL_HEPA.jpg
          Reply  Fri Apr 23 18:00:02 2021, gautam, Update, PSL, HEPA speed lowered HEPAdiag.pdf
             Reply  Fri Apr 23 19:26:58 2021, Koji, Update, PSL, HEPA speed lowered 
             Reply  Fri May 14 17:45:05 2021, rana, Update, PSL, HEPA speed raised 
                Reply  Tue May 18 00:52:38 2021, rana, Update, PSL, HEPA speed raised Untitled.png
                   Reply  Tue May 18 20:26:11 2021, rana, Update, PSL, HEPA speed raised 
    Reply  Thu Apr 22 15:15:26 2021, gautam, Update, PSL, PMC transmission 
       Reply  Thu Apr 22 15:34:54 2021, Anchal, Update, PSL, PMC transmission 
Message ID: 16144     Entry time: Tue May 18 00:52:38 2021     In reply to: 16141     Reply to this: 16145
Author: rana 
Type: Update 
Category: PSL 
Subject: HEPA speed raised 

Looks like the fan lowered the temperature as expected. Need to get a few more days of data to see if its stabilized, or if that's just a fluke.

The vertical line at 00:00 UTC May 18 is about when I turned the fans up/on.

Attachment 1: Untitled.png  50 kB  Uploaded Tue May 18 01:53:46 2021  | Hide | Hide all
Untitled.png
ELOG V3.1.3-